OpenStudy (sodapop):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2?

OpenStudy (sodapop):

\[a _{n}=5-2(n-1);all integers where n\[\ge 1\] \[a _{n}=5-2(n-1); all integers where n\[\ge 0\] \[a _{n}=5-3(n-1); all integers where n\[\ge 1\] \[a _{n}=5-3(n-1); all integers where n\[\ge 0\]

OpenStudy (sodapop):

@Hero

OpenStudy (jim766):

an = a1 + (n-1)d , a1 is the 1st term, d is the common difference . substitute in

OpenStudy (sodapop):

So the answer would be A ? @Jim766

OpenStudy (jim766):

the difference between the first term (5) and the 2nd (2) is negative 3

OpenStudy (jim766):

an = 5 + (n-1)(-3)

OpenStudy (jim766):

all they do then is move the neg 3 to the other side of the parenthese

OpenStudy (jim766):

an = 5 -3(n -1)

OpenStudy (sodapop):

Right, so how do we get the other part of the answer?

OpenStudy (jim766):

n is the number of the term you need to find... it is not possible to have a zeroth term x >= 1

OpenStudy (sodapop):

Alright cool, thank you for helping me out.

OpenStudy (anonymous):

CCCCCC is WRONG

OpenStudy (anonymous):

Who is CCCCCC? @Jasmine32012

OpenStudy (anonymous):

what was the correct answer?